Write an equation in slope-intercept form for the line that contains each pair of points below
(-9,-1) and (3,7)

Answer by rfer(16322)

m=7--1/3--9
m=8/12
m=2/3
---------------------
y=mx+b
7=2/3*3+b
7=6/3+b
5=b
----------------
y=2/3x+5